QUIZ #14: CALCULUS 1A (Stankova) Wednesday, May 5, 2004 Section 10:00–11:00 (Voight)

Name:

Please complete the following problem(s) in the space provided. You may not use a calculator. You will have 15 minutes to complete the quiz.

Please include all relevant intermediate calculations and explain your work when appropriate.

Problem 1. Find the area of the region bounded by the curves y = \sqrt{x + 1} and y = x^{2} − 7x/3 between x = 0 and x = 3.

QUIZ #14: CALCULUS 1A (Stankova) Wednesday, May 5, 2004 Section 11:00–12:00 (Voight)

Name:

Please complete the following problem(s) in the space provided. You may not use a calculator. You will have 15 minutes to complete the quiz.

Please include all relevant intermediate calculations and explain your work when appropriate.

Problem 1. Evaluate the indefinite integral

∫ 5x\sqrt{x^{2} − 1} dx.

Problem 2. Evaluate the definite integral.

∫_{1}^{4} (\frac{1}{2\sqrt{x}}) e^{\sqrt{x}} dx.